Consider a parallel plate capacitor. The magnitude of the electric field in the capacitor is E...

  1. Consider a parallel plate capacitor. The magnitude of the electric field in the capacitor is E = 2 x 103N/C, and the plate separation is 5 x 10-3m. What is the V (voltage) between the plates?
    1. 1.77 V
    2. 3.00 V
    3. 8.85 V
    4. 12.4 V
    5. 17.7 V

answer: B

  1. Continuing from the previous problem: A test charge q0= -2 x 10-3 C is released from rest at a negative plate of the capacitor. What is the kinetic energy of the test charge just before it reaches a positive plate?
    1. 6 x 10-3J
    2. 2 x 10-3J
    3. 0 J
    4. -3 x 10-3J
    5. -6 x 10-3J

answer: A

I not sure how they got those answers
Answer #1

E = 2 E+3 N/C ; d = 5.0 E-3 m

we assume the field E across the capacitor is uniform

Potential across the plates V = Ed = 2.0E+3 * 5.0E-3 = 10 V - ideal

due to fringe effect the field is not completely uniform across. at the ends of the plates the field extends slightly outside the area of the plates. Thus Ed >10 and V

V = 12.24 - (d)

test charge qo = -2E-3 C

Force on the charge F = Eqo

work done W = Fd = Eqod = KE of the charge

= 2.0E+3 * 2.0E-3 * 5.0E-3 = 20 E-3 J